OpenStudy (anonymous):

Find the indicated probability If P(B) = 0.5, P(A or B) = 0.3, and P(A and B) = 0.6. P(A) =

OpenStudy (shubhamsrg):

note that p(A & B) = p(A) + p(B) - p(A or B)

OpenStudy (dumbcow):

no i think you have it backwards P(A or B) = P(A) + P(B) - P(A and B) --> P(A) = 0.3 + 0.6 -0.5

OpenStudy (dumbcow):

these probabilities don't make sense in real world application

OpenStudy (shubhamsrg):

@dumbcow have it backwards? what do you mean.. what you wrote and what i wrote is the same thing..just a slight shuffle..

OpenStudy (dumbcow):

oh i see, you were right i never think of it that way, normally P(A or B) > P(A and B)

OpenStudy (shubhamsrg):

hmmn..